OpenStudy (sophia_onat):

Solve the following system of equations using elimination: 3x - 4y = 8 10x + 5y = 45

jhonyy9 (jhonyy9):

do you know what mean elimination method to solve a system of equations ? so in case of one equation what needed multiplie both sides by a number so that when assuming these equations these same terms eliminate so for example 3x-4y =8 in case of this equation multiplie both sides by 5 10x +5y =45 in this case multiplie both sides by 4 and so will get 15x -20y =40 40x +20y =180 ------------------ assume the same terms 55x 0 = 220 55x = 220 x = 220/55 = 4 x=4 do you can getting now easy the y value than just substitute the x value of 4 inside one equation in place of x and solve what you get for y - so 3*4 -4y = 8 12 -4y = 8 12 -8 = 4y 4 =4y divide both sides by 4 and will get y = 1 x = 4 y = 1 hope so much that this is understandably sure easy